The everyshi package^{∗}

Martin Schröder

Crüsemannallee 3 D-28213 Bremen Martin.Schroeder@ACM.org

                               PGP-Key:  2048 bit / KeyID 292814E5
                            7E86 6EC8 97FA 2995  82C3 FEA5 2719 090E

1999/06/08

Abstract

This package defines a new command \EveryShipout analogous to \AtBeginDocument etc., whose argument is executed before each \shipout.

Contents

1 Introduction 1

2 Usage 1

3 Options 2 4 Required packages 2

5 Acknowledgements 2

1 Introduction

This package provides the hook \EveryShipout whose argument is executed after the output routine has constructed \box255, and before \shipout is called.

An example application for this package would be a package for adding text to the bottom of each page. Such a package does exist: prelim2e[1].

2 Usage

\EveryShipout{⟨code⟩} declares ⟨code⟩ that is saved internally and executed before each \shipout.

The ⟨code⟩ is executed after \box255 has been constructed by the output routine and can change \box255. \shipout is called after ⟨code⟩.

Repeated use of this command is permitted: the code in the argument is stored (and executed) in the order of their declarations.

∗The version umber of this file is v2.03, last revised 1999/06/08.

The name everyshi is a tribute to the 8+3 file-naming convention of certain “operating systems”; strictly speaking it should be everyshipout.

3 Options

The package has no options.

4 Required packages

The package does not require any further packages.

5 Acknowledgements

Version 2.0 of everyshi borrows heavily from quire.tex of the Midnight Macros by Marcel R. van der Goot (marcel@cs.caltech.edu). The pointer to quire was provided by Lothar Meyer-Lerbs (TEXSatz@zfn.uni-bremen.de). As usual Rebecca Stiels improved the quality of this documentation.

References

[1] Martin Schröder. The prelim2e package.

   CTAN:tex-archive/macros/latex/contrib/supported/ms/prelim2e.dtx.

L^{A}TEX 2ε package.